Fermi National Laboratory


Accelerator Update for April 22, 2005 - April 25, 2005

Friday April 22
The day shift began with Operations monitoring store 4108, with the Antiproton Source (Pbar) stacking, with NuMI experts conducting studies, and with the MiniBooNE and Meson MCenter experiments taking beam.

At 9:38 AM, the Cryo Coordinator reported that the Tevatron (TeV) sector C2 wet engine was about to fail. Cryo system technicians were ready to put the system in JT mode.

Operations put the Recycler into shot setup to stash antiprotons at 10:08 AM. Operations completed nine transfers fifty minutes later. Any shot setup stops Pbar from stacking and halts beam to Switchyard (SY).

Pbar resumed stacking at 11:08 AM.

At 1:41 PM, cryo system techs said that the C2 wet engine wouldn't last any longer. Operations warned the experiments of the switch to JT mode, since the switch might cause the store to abort. They made the switch at 2:07 PM and began repairs on the wet engine. The techs switched back to the wet engine at 2:36 PM.

At 5:32 PM, a Booster RF station (BRF16) tripped off. It reset okay.

At 8:44 PM, BRF16 tripped off. It reset okay.

At 9:42 PM, operators began having a problem with the Booster beam, which they couldn't resolve. An expert told them to halt Booster beam and that he would come in to investigate.

Pbar resumed stacking at 11:37 PM.

Saturday April 23
The midnight shift began with Operations monitoring store 4108 and the Recycler stash, with Booster experts tuning, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Booster experts reported at 12:47 AM that they had finished their work on Booster extraction.

At 2:59 AM, the Meson beam permit tripped off due to a LCW problem in M01. This shut off MCenter beam. Operators could not reset the system and called in experts.

Operations terminated store 4108 at 4:03 AM and then put the TeV into shot setup.

A water system tech had the M01 LCW system back up by 5:01 AM. At 6:04 AM, the expert and an operators access M01 to switch the system to a different pump.

Operations established combination store 4109 at 6:26 AM with an initial luminosity of 94.5E30. Pbar resumed stacking.

At 11 AM, a Booster expert asked Operations to halt Booster beam so he could make some adjustments.

At 11:10 AM, BRF16 tripped off. It reset okay.

At 11:11 AM, the Booster expert reported that a kicker (MK05) wasn't firing properly. Operations called in a kicker expert.

At 12:48 PM, a Pbar CAMAC link went down and a permit dropped. Operators investigated and found a breaker had tripped. They reset the breaker and stacking improved at 1:23 PM.

The Recycler began stashing antiprotons at 1:36 PM. Right after the stash began, Operations prepared Booster for access to investigate the MK05 kicker.

The Recycler stash ended at 2:01 PM.

Booster experts determined at 3:34 PM that the MK05 kicker magnet would have to be replaced.

At 5:07 PM, operators accessed the Main Injector (MI) MI12B area to check on a MiniBooNE LCW system.

Booster techs reported at 5:10 PM that they had completed the MK05 kicker magnet replacement.

Pbar resumed stacking at 5:44 PM. MiniBooNE and MCenter resumed taking beam about ten minutes later.

At 7:57 PM. Operations reported that the Pbar stacking efficiency was extremely poor. Operators found a problem with the Debuncher RF station (DRF1) VME processor. Experts were called in.

Pbar experts reset a crate that cleared the DRF1 problem at 8:55 PM.

At 11 PM, the power supply (ORBMP) tripped off and wouldn't reset. Operator investigated and found a blown fuse.

Sunday April 24
The midnight shift began with Operations monitoring store 4109 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

Recycler began stashing at 12:05 AM.

At 12:58 AM, Operations reported that they were losing antiprotons out of every transfer.

The Recycler stash ended around 1:10 AM. Pbar resumed stacking and MCenter resumed taking beam about ten minutes later.

Operations terminated store 4109 at 10:01 AM.

At 10:23 AM, Operations prepared the D-Zero (D0) collision hall for an access. The access ended at 12:19 PM.

At 12:21 PM, Operations disabled the beam to MiniBooNE due to a LCW expansion tank problem. Experts were called.

At 1:44 PM, while operators were trying to tune the TeV, they ran across what seemed to be a BPM problem. They rebooted all the BPMs, but it didn't solve the problem. They contacted a TeV expert who suggested a test that showed beam was good. Operators finished their TeV tune.

A MiniBooNE expert reported at 3:10 PM that the beamline will stay off until Monday.

Operations established combination store 4111 at 5:03 PM with an initial luminosity of 55.87E30. Pbar resumed stacking. Operations reported that there was a significant loss of antiprotons in the P1 line. Operators continued to investigate.

A NuMI expert began a tune measurement study at 8:57 PM.

MCenter resumed taking beam at 8:57 PM.

Operations began preparing the Recycler to stash antiprotons at 11:14 PM.

Monday April 25
The midnight shift began with Operation monitoring store 4111, with the Recycler in shot setup, and with MiniBooNE off due a LCW problem.

The Recycler stash ended at 12:42 AM.

At 1:51 AM, an Accumulator power supply (A:LQ) tripped off, which caused the loss of the stack. Operator investigated and found the 480-volt breaker had tripped; they called in experts.

The Pbar experts reported at 6:27 AM that they had found and fixed a poor connection on A:LQ.

The Plan for Monday
The plan for today is to allow Pbar to stack to 100mA, which should occur later this afternoon, and then terminate store 4011. Operations will put the TeV back into shot setup.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/25/2005   email Fermilab